OpenStudy (anonymous):

In Problem Set 1: Problem 1J-2: How does d/dx cos(x) when x is PI/2 = Cos(x)-Cos(PI/2)/(x-PI/2). Shouldn't the numerator be Cos(x-PI/2)-Cos(PI/2)? (Difference Quotient)

OpenStudy (noelgreco):

Some people who might want to help you may not know how to find problem 1J-2. Posting a link might get you more help.

OpenStudy (anonymous):

We define derivative this way: \[ f'(x) = \lim_{\Delta x \rightarrow 0 } \frac { f(x + \Delta x) - f(x) } { \Delta x } \] Substitute \[ \Delta x = x_0 - x \] to get \[ f'(x) = \lim_{ x \rightarrow x_0 } \frac { f(x + (x_0 - x)) - f(x) } { x_0 - x } \\ = \lim_{ x \rightarrow x_0 } \frac { f(x_0) - f(x) } { x_0 - x } \\ = \lim_{ x \rightarrow x_0 } \frac { - (f(x) - f(x_0)) } { -(x - x_0) } \\ = \lim_{ x \rightarrow x_0 } \frac { f(x) - f(x_0) } { x - x_0 } \] Therefore \[ \cos'(x) = \lim_{x \rightarrow \pi/2} \frac { \cos ( x ) - \cos (\pi /2) } { x - \pi/2} = \lim_{x \rightarrow \pi/2} \frac { \cos ( x ) - \cos( \pi /2) } { x - \pi/2} \]